# G2RPO: Granular GRPO for Precise Reward in Flow Models

### v2

@@ 4.1 Singular Stochastic Sampling
Traditional flow-based GRPO methods introduce SDE sampling at every step to inject stochasticity and uniformly assign the final image reward to each step’s sampling direction. According to Eq. 4, the advantage $A_{t}^{i}$ of each sampling direction at step $t$ is equally assigned with $A_{0}^{i}$. However, the reward signal available only after multiple decision steps impedes the model’s capability to link the final reward to each decision, thereby resulting in imprecise and sparse rewards.
To acquire a dense reward for each SDE sampling direction, a simple yet effective strategy is to confine the stochasticity to the single step selected for optimization. Firstly, we designate a set of candidate SDE timesteps $M\subset\{1,\ldots,T\}$ with $|M|=K\leq T$. As shown in Fig. 2, given a prompt $\bm{c}$ and initial noise $\bm{x}_{T}$, each timestep denoted by $k\in M$ will be optimized in the training phase. Then, a common starting point $\bm{x}_{k}$ for the group is acquired using ODE sampling from Eq. 1. Notably, the Singular Stochastic Sampling strategy employs SDE sampling only at $\bm{x}_{k}$ and samples $G$ distinct denoising directions to get the next noisy state $\{\bm{x}_{k-1}^{i}\}^{G}_{i=1}$. Then each $\bm{x}_{k-1}^{i}$ undergoes k1k-1 steps of ODE sampling to generate a deterministic denoised image $\bm{x}_{0\leftarrow k}^{i}$. Based on our sampling strategy, the variance of the group reward $\{R(\bm{x}_{0\leftarrow k}^{i},\bm{c})\}^{G}_{i=1}$ is entirely determined by the distinct denoising directions introduced by the SDE sampling at step $k$. And a step-aware, precise advantage can be acquired:
$A_{k}^{i}=\frac{R(\bm{x}_{0\leftarrow k}^{i},\bm{c})-\operatorname{mean}(\{R(\bm{x}_{0\leftarrow k}^{i},\bm{c})\}_{i=1}^{G})}{\operatorname{std}(\{R(\bm{x}_{0\leftarrow k}^{i},\bm{c})\}_{i=1}^{G})}.$ (8)
@@
$f(r,A,\theta,\varepsilon,\beta)=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{K}\sum_{k\in M}\left(\min\left(r_{k}^{i}(\theta)A_{k}^{i},\operatorname{clip}\left(r_{k}^{i}(\theta),1-\varepsilon,1+\varepsilon\right)A_{k}^{i}\right)\right).$ (9)
After the sampling phase is completed, the training of GRPO requires the computation of $p_{\theta}\left(\bm{x}_{k-1}^{i}\mid\bm{x}_{k}^{i},\bm{c}\right)$ to obtain $r_{k}^{i}(\theta)$ refer to eq 7. In practice, each distinct sampling starting point $\bm{x}_{k}^{i}$ needs to be fed into the flow model to compute the corresponding ODE denoising direction $\bm{v}_{k}^{i}$. However, our sampling strategy shares a common starting point $\bm{x}_{k}$, allowing a group of $G$ samples to reuse the same $\bm{v}_{k}$, which in turn improves training efficiency.
@@ 4.2 Multi-Granularity Advantage Integration
Singular Stochastic Sampling accurately constrains stochasticity into the single SDE step, ensuring a strong correlation between the reward and the injected noise. Nevertheless, how to acquire a comprehensive reward for the denoising direction of the current step still requires further investigation.
As shown in Fig. 3, we observe that under identical $\bm{x}_{k}$ and prompt $\bm{c}$ conditions, the denoising trajectory generated by singular stochastic sampling is not robust when assessing the corresponding SDE denoising direction. Images generated from denoising trajectories with different granularities show similar overall content but exhibit discrepancies in detail due to the varying denoising intervals. Such differences are evident in the scores assigned by the Reward Model, further influencing the numerical values of the advantage within the group, and even the optimization direction.
To this end, we propose a Multi-Granularity Advantage Integration module to perform multi-granularity denoising on the sampled denoising directions within a group. The advantages of the images denoised at different granularities are then integrated to form the final evaluation. Specifically, as shown in Fig. 2, step $k$ is the SDE sampling step, and $G$ distinct denoising directions are sampled to acquire next noisy state $\{\bm{x}_{k-1}^{i}\}^{G}_{i=1}$. Under the conventional granularity condition, each $\bm{x}_{k-1}^{i}$ undergoes k1k-1 steps to obtain the final denoised image. The sequence of denoising timesteps can be represented as: $\mathcal{S}=\{1,2,\ldots,k-1\}$. For our Multi-Granularity denoising module, a set of integer scaling factors $\Lambda=\{\lambda_{1},\lambda_{2},...,\lambda_{j}\},|\Lambda|=J$ is defined to represent different denoising granularities. Each $\lambda_{j}$ implements interval sampling for different denoising granularity, that means sample every $\lambda_{j}$-th step from the total k1k-1 steps. The denoising timestep sequence $\mathcal{S}_{j}$ can be formally represented as:
@@
After $N_{j}$ subsequent steps denoising for $\{\bm{x}_{k-1}^{i}\}^{G}_{i=1}$, a group of noise-free images $\{x_{0}^{i,j}\}^{G}_{i=1}$ are generated. Subsequently, different groups images gets the reward $\{R(\bm{x}_{0\leftarrow k}^{i,j},\bm{c})\}^{G}_{i=1}$ from a reward model and then compute the intra-group advantages $\{A_{t}^{i,j}\}^{G}_{i=1}$ with Eq. 4. Similar to the joint training with multiple reward models (e.g., HPS-v2.1 and CLIP Score) in DanceGRPO, where the advantages from different reward models are directly summed, we combines the advantages from different granularities to get $\{A_{t}^{i,\text{mix}}\}^{G}_{i=1}$:
@@
Finally, $f(r,A,\theta,\varepsilon,\beta)$ is updated to:
$f(r,A,\theta,\varepsilon,\beta)=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{K}\sum_{k\in M}\left(\min\left(r_{k}^{i}(\theta)A_{k}^{i,\text{mix}},\operatorname{clip}\left(r_{k}^{i}(\theta),1-\varepsilon,1+\varepsilon\right)A_{k}^{i,\text{mix}}\right)\right),$ (12)
and using Eq. 5 to optimize the policy $\pi_{\theta}$. A detailed algorithm is illustrated in Algorithm 1.
Algorithm 1 G2RPO Training Process
@@
7:   Init same noise $\bm{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$
@@
13:      SDE Sampling a group samples: $\bm{x}_{k-1}^{i}$  // $i$-th direction in the group
14:     else if $t>k$ then
@@
16:       ODE Sampling with granularity $\lambda_{j}$: $\bm{x}_{t-1}^{i,j}$
@@
20:    Get a group of reward: $R(\bm{x}_{0\leftarrow k}^{i,j})$
21:    $A_{k}^{j}\leftarrow\frac{R(\bm{x}_{0\leftarrow k}^{i,j},\bm{c})-\mu^{j}}{\sigma^{j}}$
@@ 5.1 Implementation Details
Datasets and Backbone. Following DanceGRPO (Xue et al., 2025) and MixGRPO (Li et al., 2025), we evaluate our G2RPO using the HPSv2 (Wu et al., 2023) dataset. It contains 103,700 text prompts for training and 400 diverse prompts for testing. The text-to-image model employed for reinforcement learning is Flux.1-dev (Labs, 2024), a leading flow model in the community.
@@
Evaluation Metrics. To evaluate the effectiveness and robustness of our G2RPO, multiple reward models are applied as evaluation metrics. These reward models assess the alignment between generated images and human preferences from multiple dimensions. Specifically, HPS-v2.1 (Wu et al., 2023), CLIP Score (Radford et al., 2021), and Pick Score (Kirstain et al., 2023) collectively assess semantic alignment and visual coherence. Image Reward (Xu et al., 2023) focuses on visual quality and aesthetic appeal, while Unified Reward (Wang et al., 2025) is the SOTA unified reward model that comprehensively evaluates both alignment with the caption and overall image quality.
@@
Evaluation Setting. Similar to DanceGRPO and MixGRPO, two experimental settings are employed. Firstly, a single HPS-v2.1 reward model is utilized for training to verify the upper limit of improvement for in-domain performance. However, as demonstrated by DanceGRPO, HPS-v2.1 is prone to model hacking due to biases in the training set, leading to degradation in other evaluation metrics. Therefore, our primary experiment also involves joint training with both HPS-v2.1 and CLIP Score as reward models to acquire stable and robust results.
@@
Sampling Phase. Following DanceGRPO, a shared initialization noise is used to generate a group of $12$ images from the same text prompt. The total sampling step $T=16$ to enhance computational efficiency and the advantage clip $\varepsilon=5$ in Eq. 6. Note that, the parameter $\eta$ in Eq. 3 directly determines the noise level, which in turn defines the size of the stochastic exploration space in the SDE. Leveraging the Singular Stochastic Sampling strategy, our precise reward can tolerate a larger $\eta=0.7$. The set of candidate SDE timesteps $M$ consists of the first 8 timesteps to improve training efficiency. For the Multi-Granularity Advantage Integration strategy, the set of distinct granularities $\Lambda=\{1,2,3\}$.
@@
Training Phase. All experiments are conducted using $16\times$ NVIDIA H200 GPUs, with a batch size of $1$. The AdamW optimizer is used, configured with a learning rate of $2\times 10^{-6}$ and a weight decay of $1\times 10^{-4}$. Mixed precision training is implemented using bfloat16 (bf16) format. The training iteration is $300$.
5.2 Main Results
Quantitative Evaluation. As shown in Tab. 1, DanceGRPO and MixGRPO serve as the baselines for comparison with our G2RPO. Additionally, the performance of employing the Singular Stochastic Sampling strategy without multi-granularity denoising (G2RPO w/o MGAI) is also proposed. It can be observed that when HPS-v2.1 is used solely as the training reward model, our Singular Stochastic Sampling achieves a relative improvement of 6.52% compared to DanceGRPO. This indicates that constraining the stochasticity to a single step yields a precise reward signal, which in turn provides a more faithful optimization signal and enables GRPO to enhance the optimization ceiling. However, as demonstrated by DanceGRPO, optimizing solely with HPS-v2.1 can induce model hacking, which in turn compromises other out-of-domain evaluation metrics. Secondly, for the setting of multi-reward (HPS-v2.1 and CLIP Score) optimization, the results indicate that G2RPO attains superior performance across both in-domain and out-of-domain rewards. Under the multi-granularity denoising condition, groups of images at different granularities provide a more comprehensive evaluation for the SDE sampling direction. This multi-granularity paradigm also allows for more flexible adaptation to the preferences of various reward models, thereby achieving significant improvements in various out-of-domain dimensions.
Table 1: Quantitative Results. Comparison of results on in-domain and out-of-domain rewards.
Reward Model Method In-Domain Out-of-Domain
HPS-v2.1 CLIP Score Pick Score ImageReward Unified Reward
@@ 5.2 Main Results
Qualitative Comparison. Fig. 4 presents the qualitative comparison among the original Flux.1-dev, DanceGRPO, MixGRPO, and our proposed G2RPO. It can be observed that G2RPO provides enhanced detail fidelity and improves the consistency with the text prompt, achieving superior alignment with human preferences. For instance, in the second column, our G2RPO faithfully captures the specified expressions and even the nuances of the chess pieces as described in the prompt, delivering finer details and higher visual quality. Moreover, in the “poster” case depicted in the last column, G2RPO not only adheres to the spatial requirement of a clear left-right demarcation but also renders the reflections of the trees with remarkable clarity. Additionally, the overall style of the image generated by G2RPO is more consistent with the aesthetic demands of poster design.
5.3 Ablation Study
@@
Reward Model $\Lambda$ In-Domain Out-of-Domain
HPS-v2.1 CLIP Score Pick Score ImageReward Unified Reward
HPS-v2.1 & CLIP $\{1\}$ 0.372 0.395 0.234 1.421 3.688
@@
5.4 Further exploration of varying inference steps
In the main experiment Tab. 1, we maintained the same inference settings as DanceGRPO and MixGRPO, generating images with a resolution of $1024\times 1024$ in $50$ steps. Notably, our MGAI module, which evaluates denoising samples of different granularities in a mixed manner, provides a more comprehensive assessment. With the assistance of this module, G2RPO exhibits stronger robustness to varying denoising step configurations. As shown in Tab. 3, all flow-based GRPO methods are jointly trained with HPS-v2.1 and CLIP. When the total inference timesteps are reduced to $20$ or even $10$ steps, G2RPO still achieved significant performance improvements across various in-domain and out-of-domain evaluation reward models.

### v3

@@ 3.2 Singular Stochastic Sampling
Traditional flow-based GRPO methods introduce ODE-to-SDE conversion sampling at every step to inject stochasticity and uniformly assign the final image reward to each step’s sampling direction. According to Eq. 4, the advantage $A_{t}^{i}$ of each sampling direction at step $t$ is equally assigned with $A_{0}^{i}$. Notably, the reward signal available only after multiple decision steps impedes the model’s capability to link the final reward to each decision, thereby resulting in imprecise and sparse rewards.
In accordance with the design of DanceGRPO, we define $M$ as the first half of the denoising timesteps, i.e., $M=\{T,T-1,\ldots,\lfloor T/2\rfloor\}$. This selection prioritizes the early denoising steps, where the exploration space of the SDE is larger and significantly determines the overall direction of the entire denoising chain. In contrast, the improvements from GRPO in the later stages are relatively minor. Consequently, training is conducted exclusively on these timesteps after each round of group image sampling, thereby optimizing computational efficiency. Moreover, to acquire a dense reward for each SDE sampling direction, a simple yet effective strategy is to confine the stochasticity to the single step selected for optimization. As shown in Fig. 2, given a prompt $\boldsymbol{c}$ and initial noise $\boldsymbol{x}_{T}$, for each timestep $k\in M$, a common starting point $\boldsymbol{x}_{k}$ for the group is acquired using ODE sampling from Eq. 1. Then, our Singular Stochastic Sampling strategy employs SDE sampling only at $\boldsymbol{x}_{k}$ and samples $G$ distinct denoising directions to get the next noisy state $\{\boldsymbol{x}_{k-1}^{i}\}^{G}_{i=1}$. Each $\boldsymbol{x}_{k-1}^{i}$ undergoes k1k-1 steps of ODE sampling to generate a deterministic denoised image $\boldsymbol{x}_{0\leftarrow k}^{i}$. Based on our sampling strategy, the variance of the group reward $\{R(\boldsymbol{x}_{0\leftarrow k}^{i},\boldsymbol{c})\}^{G}_{i=1}$ is entirely determined by the distinct denoising directions introduced by the SDE sampling at step $k$. A step-aware, precise advantage can be acquired:
$A_{k}^{i}=\frac{R(\boldsymbol{x}_{0\leftarrow k}^{i},\boldsymbol{c})-\operatorname{mean}(\{R(\boldsymbol{x}_{0\leftarrow k}^{i},\boldsymbol{c})\}_{i=1}^{G})}{\operatorname{std}(\{R(\boldsymbol{x}_{0\leftarrow k}^{i},\boldsymbol{c})\}_{i=1}^{G})}.$ (8)
@@
$f(r,A,\theta,\varepsilon,\beta)=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{K}\sum_{k\in M}\biggl(\min\Bigl(r_{k}^{i}(\theta)A_{k}^{i},$
$\operatorname{clip}\bigl(r_{k}^{i}(\theta),1-\varepsilon,1+\varepsilon\bigr)A_{k}^{i}\Bigr)\biggr).$ (9)
After the sampling phase is completed, the training of GRPO requires the computation of $p_{\theta}\left(\boldsymbol{x}_{k-1}^{i}\mid\boldsymbol{x}_{k}^{i},\boldsymbol{c}\right)$ to obtain $r_{k}^{i}(\theta)$ refer to eq 7. In practice, each distinct sampling starting point $\boldsymbol{x}_{k}^{i}$ needs to be fed into the flow model to compute the corresponding ODE denoising direction $\boldsymbol{v}_{k}^{i}$. However, our sampling strategy shares a common starting point $\boldsymbol{x}_{k}$, allowing a group of $G$ samples to reuse the same $\boldsymbol{v}_{k}$, which in turn improves training efficiency.
@@ 3.3 Multi-Granularity Advantage Integration
Singular Stochastic Sampling accurately constrains stochasticity into the single SDE step, ensuring a strong correlation between the reward and the injected noise. Then, how to acquire a comprehensive reward signal for the denoising direction of the current step still requires further investigation.
As shown in Fig. 3, we observe that under identical $\boldsymbol{x}_{k}$ and prompt $\boldsymbol{c}$ conditions, the denoising trajectory generated by singular stochastic sampling is not robust when assessing the corresponding SDE denoising direction. Specifically, while the overall content of the images generated from different denoising trajectories remains similar, the details vary significantly due to the differing denoising intervals. This variation in denoising granularity leads to discrepancies in the scores assigned by the reward model, which in turn influences the numerical values of the advantage within the group and even the optimization direction.
To conduct a comprehensive assessment of the single-step SDE sampling direction, we propose a Multi-Granularity Advantage Integration module, which performs multi-granularity denoising on the sampled denoising directions within a group. The advantages of the images denoised at different granularities are then integrated to form the final evaluation. Specifically, as shown in Fig. 2, step $k$ is the SDE sampling step, and $G$ distinct denoising directions are sampled to acquire next noisy state $\{\boldsymbol{x}_{k-1}^{i}\}^{G}_{i=1}$. Under the conventional granularity condition, each $\boldsymbol{x}_{k-1}^{i}$ undergoes k1k-1 steps to obtain the final denoised image. The sequence of denoising timesteps can be represented as: $\mathcal{S}=\{1,2,\ldots,k-1\}$. For our Multi-Granularity denoising module, a set of integer scaling factors $\Lambda=\{\lambda_{1},\lambda_{2},...,\lambda_{j}\},|\Lambda|=J$ is defined to represent different denoising granularities. Each $\lambda_{j}$ implements interval sampling for different denoising granularity, that means sample every $\lambda_{j}$-th step from the total k1k-1 steps. The denoising timestep sequence $\mathcal{S}_{j}$ can be formally represented as:
@@
After $N_{j}$ subsequent steps denoising for $\{\boldsymbol{x}_{k-1}^{i}\}^{G}_{i=1}$, a group of noise-free images $\{x_{0}^{i,j}\}^{G}_{i=1}$ are generated. Subsequently, different groups images gets the reward $\{R(\boldsymbol{x}_{0\leftarrow k}^{i,j},\boldsymbol{c})\}^{G}_{i=1}$ from a reward model and then compute the intra-group advantages $\{A_{t}^{i,j}\}^{G}_{i=1}$ with Eq. 4. Similar to the joint training with multiple reward models (e.g., HPS-v2.1 and CLIP Score) in DanceGRPO, where the advantages from different reward models are directly summed to provide a multi-dimensional comprehensive evaluation of a single denoising direction, we combines the advantages from different granularities to get $\{A_{t}^{i,\text{mix}}\}^{G}_{i=1}$:
@@
Accordingly, $f(r,A,\theta,\varepsilon,\beta)$ in Eq 3.2 can be updated to the following formula:
$f(r,A,\theta,\varepsilon,\beta)=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{K}\sum_{k\in M}(\min(r_{k}^{i}(\theta)A_{k}^{i,\text{mix}},$
$\operatorname{clip}\left(r_{k}^{i}(\theta),1-\varepsilon,1+\varepsilon\right)A_{k}^{i,\text{mix}})),$ (12)
Finally, our G2RPO uses Eq. 5 to optimize the policy $\pi_{\theta}$ across all timesteps in the set $M$. A detailed algorithm is illustrated in Algorithm 1.
Algorithm 1 Granular-GRPO Training Process
@@
7:   Init same noise $\boldsymbol{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$
@@
13:      SDE Sampling a group samples: $\boldsymbol{x}_{k-1}^{i}$
14:     else if $t<k$ then
@@
16:       ODE Sampling with granularity $\lambda_{j}$: $\boldsymbol{x}_{t-1}^{i,j}$
@@
20:    Get a group of reward: $R(\boldsymbol{x}_{0\leftarrow k}^{i,j})$
21:    $A_{k}^{j}\leftarrow\frac{R(\boldsymbol{x}_{0\leftarrow k}^{i,j},\boldsymbol{c})-\mu^{j}}{\sigma^{j}}$
@@ 4.1 Implementation Details
Datasets and Backbone. Following the same setting in DanceGRPO [xue2025dancegrpo] and MixGRPO [li2025mixgrpo], the prompt dataset HPSv2 [wu2023human] is utilized. It contains 103,700 text prompts for training and 400 diverse prompts for testing. The text-to-image model employed for reinforcement learning is Flux.1-dev [flux2024], a leading flow model in the community.
@@
Evaluation Metrics. To evaluate the effectiveness and robustness of our G2RPO, multiple reward models are applied as evaluation metrics. These reward models assess the alignment between generated images and human preferences from multiple dimensions. Specifically, HPS-v2.1 (HPS) [wu2023human], CLIP Score (CLIP) [radford2021learning], and Pick Score (PS) [kirstain2023pick] collectively assess semantic alignment and visual coherence. Image Reward (IR) [xu2023imagereward] focuses on visual quality and aesthetic appeal, while Unified Reward (UR) [wang2025unified] is the SOTA reward model that leverages the powerful visual understanding capabilities of LLM comprehensively evaluates both alignment with the caption and overall quality.
@@
Evaluation Setting. Similar to DanceGRPO and MixGRPO, two experimental settings are employed. Firstly, a single HPS-v2.1 reward model is utilized for training to verify the upper limit of improvement for in-domain performance, i.e., both training and testing are conducted using the same reward model. However, as demonstrated by DanceGRPO, HPS-v2.1 is prone to model hacking due to biases in the training set, leading to degradation in other evaluation metrics. Similarly, our primary experiment involves joint training with both HPS-v2.1 and CLIP Score as reward models to acquire stable and robust results. Subsequently, we evaluate the model using out-of-domain metrics, i.e., reward models that were not seen during training.
@@
Sampling Phase. Following DanceGRPO, a shared initialization noise is used to generate a group of $12$ images from the same text prompt. The total sampling step $T=16$ to enhance computational efficiency and the advantage clip $\varepsilon=5$ in Eq. 3.1. The parameter $\eta=0.7$ in Eq. 3.1 directly determines the noise level, which in turn defines the size of the stochastic exploration space in the SDE. All comparison methods share the aforementioned parameter configuration. For the Multi-Granularity Advantage Integration strategy, the set of distinct granularities $\Lambda=\{1,2,3\}$.
@@
Training Phase. All experiments are conducted using $16\times$ NVIDIA H200 GPUs, with a batch size of $1$. The AdamW optimizer is used, configured with a learning rate of $2\times 10^{-6}$ and a weight decay of $1\times 10^{-4}$. Mixed precision training is implemented using bfloat16 (bf16) format.
Table 1: Quantitative Results. Comparison of results on various reward metrics.
Reward Model Method HPS CLIP PS IR UR
@@
4.2 Main Results
Quantitative Evaluation. As shown in Tab. 1, DanceGRPO and MixGRPO serve as the baselines for comparison with our G2RPO. Flow-GRPO is not included since it employs a, similar ODE-to-SDE conversion across all timesteps as contemporary DanceGRPO, but fit a specific reward target. In addition, the Singular Stochastic Sampling strategy without multi-granularity denoising (G2RPO w/o MGAI) is also proposed. It can be observed that when HPS-v2.1 is used solely as the training reward model, our Singular Stochastic Sampling achieves a relative improvement of 6.52% compared to DanceGRPO. This indicates that constraining the stochasticity to a single step yields a precise reward signal, which in turn provides a more faithful optimization signal and enables GRPO to enhance the optimization ceiling. However, as demonstrated by DanceGRPO, optimizing solely with HPS-v2.1 can induce model hacking, which in turn compromises other out-of-domain evaluation metrics. Secondly, for the setting of multi-reward (HPS-v2.1 and CLIP Score) optimization, the results indicate that G2RPO outperforms other baselines across both in-domain and out-of-domain rewards. Under the multi-granularity denoising condition, groups of images at different granularities provide a more comprehensive evaluation of the SDE sampling direction. This robust assessment facilitates improvements in out-of-domain metrics, particularly on Unified Reward enhanced with LLM knowledge.
@@ 4.2 Main Results
Metrics $\Lambda$ HPS CLIP PS IR UR
HPS
& CLIP
$\{1\}$ 0.372 0.395 0.234 1.421 3.688
@@
Qualitative Comparison. Fig. 4 presents the qualitative comparison among the original Flux.1-dev, DanceGRPO, MixGRPO, and our proposed G2RPO. It can be observed that G2RPO provides enhanced detail fidelity and improves the consistency with the text prompt, achieving superior alignment with human preferences. For instance, in the second column, our G2RPO faithfully captures the specified expressions and even the nuances of the chess pieces as described in the prompt, delivering finer details and higher visual quality. Moreover, in the “poster” case depicted in the last column, G2RPO not only adheres to the spatial requirement of a clear left-right demarcation but also renders the reflections of the trees with remarkable clarity. The overall style of the image generated by G2RPO is more consistent with the aesthetic demands of poster design.
4.3 Ablation Study